OpenStudy (anonymous):

Solve for x: 5x+2x-4=4x+5

OpenStudy (anonymous):

X=3 right?

OpenStudy (anonymous):

5x+2x=7x. So, this is pretty much 7x-4=4x+5. Subtracting 4x from both sides gives you 3x-4=5. Adding 4 to both sides gives you 3x=9. Divide both sides by 3, x=3.